6. The electric potential in a region of space is given by V-2xy + 3уг in units of V. The electric field, in V/m, in this region is A) 2yi+(2x +6y) j B) 2y(i+3) c) -2y(i+3j) E) none of the above

I'm not sure what this question is asking. Please explain and show step by step. Thank you! :)
Answer #1

6.  

Result : (D) -2y \hat{i} - (2x+6y) \hat{j}

Explanation:

Known V = 2xy + 3y2

The electric field is given by:

\vec{E} =- \bigtriangledownV = -[\hat{i} \frac{\partial}{\partial x} + \hat{j} \frac{\partial}{\partial y} + \hat{k} \frac{\partial}{\partial z} ]V

= -[\hat{i} \frac{\partial}{\partial x} + \hat{j} \frac{\partial}{\partial y} + \hat{k} \frac{\partial}{\partial z} ] [2xy + 3y2]

= - [(2y)\hat{i} +(2x+6y)\hat{j} ]

= -(2y)\hat{i} - (2x+6y)\hat{j}